OpenStudy (anonymous):

can someone give me another easy way to figure out answers if you have alot of decimals i already tryed the whole multiplay by the highest place valuse it confuses me more and i get huge numbers !!! .02(6t-3)=.12(t-2)+.18

OpenStudy (anonymous):

\[\frac{2}{100} (6 t-3)= \frac{12}{100} (t-2)+\frac{18}{100}\]

OpenStudy (anonymous):

\[-\frac{3}{50}+\frac{3 t}{25}=-\frac{3}{50}+\frac{3 t}{25}\]The LHS equal the RHS. Set t to anything the the equation will be true.